Show all the solutions. Round off answers to two decimal places. Be consistent with the units

2. a)  Considering the two wires of the same material and cross-sectional area, how does the resistance vary with the length?

  b)  Considering the two wires of the same material and length, how does the resistance vary with the cross-sectional area?




1
Expert's answer
2021-04-20T16:49:02-0400

Explanations & calculations


  • To analyse the behaviour in each case, the equation to be used is "\small R =\Large\frac{\rho\cdot L}{A}" .
  • For the case number 1, the material & the cross sectional area are identical for both wires.
  • Since material is the same, the quality dependant on material: "\rho" is the same.
  • Then we can re-write the equation in proportional form

"\qquad\qquad
\begin{aligned}
\small R&=\small (\frac{\rho}{A})\cdot L\\
&\propto kL
\end{aligned}"

  • Then it is obvious that the resistance then changes linearly with the length of the wire such that as the length increases, the resistance increases.


2.

  • For this case, length & the resistivity: "\rho" are identical for both wires.
  • Then performing a similar analysis yeilds the following,

"\qquad\qquad
\begin{aligned}
\small R&= \small \frac{(\rho L)}{A}\\
&=\small \frac{k_1}{A}\\
&\propto\frac{1}{A}
\end{aligned}"

  • Then it could be seen that the resistance is inversly proportional to the cross sectional area such that as the cross sectional area is increased, the resistance is decreased.